OpenStudy (anonymous):

You are ordering a meal and have $15 to spend. The restaurant charges 6% sales tax and you want to leave a 15% tip. c=x+0.06x+0.15x

OpenStudy (anonymous):

what exactly is the question?

OpenStudy (anonymous):

How much can I spend for the meal?

OpenStudy (anonymous):

15 - 15*(0.06+0.15) = amount

OpenStudy (anonymous):

Or if you want it as a formula x - x(0.21) = c

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

np

OpenStudy (jdoe0001):

You are ordering a meal and have $15 to spend. The restaurant charges 6% sales tax and you want to leave a 15% tip. \( {\color{brown}{ c=x+0.06x+0.15x }} \) wondering what that was doing there

OpenStudy (anonymous):

dunno, but they would be paying you for eating there if they used that formula c = x + 0.06*x + 0.15*x haha